16 It is known that computer system programs use 32-bit virtual addresses to access storage units....

16

It is known that computer system programs use 32-bit virtual addresses to access storage units. If the physical memory space of the computer system is 1GB, and the paging management mechanism is adopted, the page size is 4KB, and each page table entry is 4B. If only one level of page table is used to realize the mapping from virtual address to physical address, how much memory space does the page table occupy?

  • A.

    1MB

  • B.

    4KB

  • C.

    1KB

  • D.

    4MB

17

What is the meaning of cache direct mapping strategy in computer storage system?

  • A.

    A line of information in memory can be mapped to any line in the cache

  • B.

    A line of information in the cache only be mapped to a specific line in memory

  • C.

    A line of information in memory can be mapped to some specific lines in the cache

  • D.

    A line of information in memory can only be mapped to a specific line in the cache

18

When the computer cache uses a direct mapping strategy, which of the following is the correct statement about the cache controller processing the physical address given by the CPU?

  • A.

    The physical memory address is divided into three parts: the lower part is used as the byte offset in the cache line, the middle part is used as the way index of the cache, and the remaining high part is used as the cache line tag

  • B.

    The physical memory address is divided into three parts: the lower part is used as the byte offset in the cache line, the middle part is used as the cache line index, and the remaining high part is used as the cache line tag

  • C.

    The physical memory address is divided into three parts: the lower part is used as the byte offset in the cache line, the middle part is used as the group index of the cache, and the remaining high part is used as the cache line tag

  • D.

    The physical memory address is divided into two parts: the lower part is used as the byte offset in the cache line, and the remaining high part is used as the cache line tag

19

It is known that the physical storage space of a computer system is 4GB, the cache data capacity is 2MB, each line can store 128B data, and the direct mapping strategy is adopted. How much extra storage space needs to be equipped with the cache controller to store other information?

  • A.

    192KB

  • B.

    16KB

  • C.

    16Kb

  • D.

    192Kb

20

It is known that the physical storage space of a computer system is 4GB, the cache data capacity is 2MB, each line can store 128B data, and the fully associative mapping strategy is adopted. How much extra storage space needs to be equipped with the cache controller to store other information?

  • A.

    416Kb

  • B.

    416KB

  • C.

    192Kb

  • D.

    192KB
Answer #1

16.

Total # of pages in virtual page = 232/212 = 1M

Size of each entry of page table = 4B

# of entries in page table = # of pages in virtual space.

Total size of page table = 1M * 4B = 4MB

Correct option is (d)

17.

Option (D) is correct option.

In direct mapped caching, a line information of main memory can be mapped to only one specific line of cache.

18.

Option (B) is the correct option.

Tag bits index bits block offset

19.

Total # of block inside cache = 2MB/128B = 16384 = 2^14

Block offset = log 128 = 7 bits.

Tag = 32-(7+14) = 11 bits.

Each cache block will have 11 bits tag + 1 bit valid bit.

So total storage bits = 12* # of cache block = 12* 16K = 192K b

Option (d) is the correct option.

20.

In fully associative cache, #of index bits = 0

Therefore tag bits = 32 - 7 = 25 bits.

One valid bit.

So total # of bits = 25+1 = 26 => 26*16K = 416Kb

Option (A) is the correct option..

Please don't post so many questions at once. Also if you have any questions comment down and please? upvote thanks...